# Nonlinear Ringdowns as Sources and Detectors of Quantum Gravitational Waves

### v1

@@ VI Ringdown nonlinearities as detectors
In quantum optics, it has long been recognized that the output of a nonlinear process strongly depends on the particle number statistics of the input parent modes [37, 36, 38]. This effect can be understood through simple physical reasoning. Particle number statistics determines whether the parent fields are bunched or anti-bunched [67], and since the occurrence of nonlinear interactions requires two or more particles, bunched states are converted more efficiently than anti-bunched ones. This is true of nonlinear processes in general, beyond SHG [40]. By analogy to quantum optics, we can therefore anticipate that the properties of nonlinearly generated GWs will depend on the graviton number statistics and other higher-order moments of the parent modes. Given the strongly nonlinear character of gravity, these might be significant effects in ringdowns.
@@
To quantify how strong nonlinear GW couplings in ringdowns can be, we may estimate the ‘graviton conversion efficiency’ between modes, compared to other types of graviton detectors. The mean number of gravitons in a GW of frequency $\omega$ and amplitude $A_{\omega}$ is [13],
$\langle N_{\omega}\rangle\approx\frac{V}{32\pi G}\omega A_{\omega}^{2}$ (23)
@@
$\eta_{\mathrm{ringdown}}=\frac{\langle N_{2\omega}\rangle}{\langle N_{\omega}\rangle}\approx 2\mathcal{R}A_{2\omega}$ (24)
where we recall the gravitational wave amplitude $A_{2\omega}\equiv A_{(4,4)}^{(2,2,0)\times(2,2,0)}$. At the source, GW amplitudes can be of order unity [80]; taking $A^{(2,2,0)\times(2,2,0)}_{(4,4)}\sim 0.1$ and $\mathcal{R}\sim 0.15$ [28] we have the nonlinear ringdown conversion efficiency
$\eta_{\mathrm{ringdown}}\sim 10^{-2}\ .$ (25)
@@
$\eta_{\mathrm{gertsenshtein}}\sim 10^{-33}$ (26)
while for the bar detector proposed in [18],
$\eta_{\mathrm{bar}}\sim 10^{-36}$ (27)
This comparison shows that ringing BHs are essentially strongly coupled nonlinear bar detectors, with a high conversion efficiency for gravitons in different modes. Nonlinear ringdowns could serve as efficient detectors of quantum properties of GWs.
Let us see an example illustrating how the statistical properties of parent states can become imprinted upon nonlinear modes in the case of SHG. Consider the dynamical eqs. (5) and (6). We take mode $a$ to be initially populated by some quantum state $\rho_{a}$, while $b$ is initially in the vacuum state. For short times, the mean number of particles in mode $b$ grows as [37],
$\langle N_{b}(t)\rangle=g^{(2)}(0)\ \langle N_{a}(0)\rangle^{2}\ \kappa^{2}t^{2},$ (28)
@@
$g^{(2)}(0)=\frac{\langle a^{\dagger}a^{\dagger}aa\rangle}{\langle a^{\dagger}a\rangle^{2}}$ (29)
is the second-order field auto-correlation function and expectation values are taken with respect to the initial state $\rho_{a}$. In deriving (28) we assumed $\sqrt{2g^{(2)}(0)\langle N_{a}(0)\rangle}\ \kappa t\ll 1$ and $t\ll\gamma^{-1}$, but note $\langle N_{b}(t)\rangle$ also exhibits significant dependence on the parent quantum state beyond the short time or small damping approximation [37].
The result (28) tells us that for short times, the efficiency with which particles are converted from modes $a$ to $b$ is directly proportional to the field auto-correlation function (29), which is related to the bunching or anti-bunching character of the wave [67]. For coherent states $g^{(2)}(0)=1$, for thermal states $g^{(2)}(0)=2$ and squeezed states have $g^{(2)}(0)=3+\frac{1}{\langle N_{a}(0)\rangle}$. Importantly, for large mean number of gravitons, squeezed states have $g^{(2)}(0)\approx 3$ independent of $\langle N_{a}(0)\rangle$. Therefore, we can expect second harmonic GW generation in BH ringdowns to be significantly enhanced for parent thermal and squeezed states in comparison to coherent states. Interestingly, second harmonic also generates squeezing in the parent mode, as shown by Mandel [65].
@@
$\langle Q_{b}(t)\rangle\approx\frac{\kappa t}{2}\Bigl{<}Q_{a}(0)P_{a}(0)+P_{a}(0)Q_{a}(0)\Bigr{>}$ (30)
which to leading order remains valid in the presence of dissipation; see the Appendix for details.
@@
$\frac{1}{2}\Bigl{<}Q_{a}P_{a}+P_{a}Q_{a}\Bigr{>}=\langle Q_{a}\rangle\langle P_{a}\rangle$
$+(2\bar{n}+1)\sinh r\sin\phi\cos\phi$ (31)
The first term on the r.h.s. is due to the coherent displacement of the GW and is the only term present for coherent states. In contrast, the second term depends on non-coherent properties of the state, the mean thermal occupation number $\bar{n}$ and squeezing parameter $\xi$. Note the non-coherent term is accompanied by an exponential enhancement characteristic of squeezing [13, 16, 17, 14, 83, 84].
As we have seen, interferometric detectors are sensitive to the mean GW quadratures [see (22)]. Eq. (30) is interesting, as it shows that SHG encodes information on the covariance of $a$ in the mean quadrature of $b$, which can be read-out by interferometric detectors. Detection of $\langle Q_{b}(t)\rangle$ by standard GW observations could therefore signal deviations of mode $a$ from a coherent state.
VII Discussion
In conclusion, SHG has been observed in numerical simulations of nonlinear BH ringdowns [28, 29]. In a quantum mechanics, second harmonics can produce non-classical states both in the parent and nonlinear modes, starting from initial coherent states [65]. This can be understood in terms of particle creation and annihilation: nonlinearities temper with particle number distributions, leading to states which classical theory cannot account for [25].
@@ Appendix A Energy conservation and frequency doubling
$[a,a^{\dagger}]=[b^{\dagger},b]=\mathds{1}$ (32)
$[a,b]=[a^{\dagger},b]=[a,b^{\dagger}]=[a^{\dagger},b^{\dagger}]=0$ (33)
@@
$J=N_{a}+2N_{b}$ (34)
For convenience we list all the commutators with third-order operators:
$[a,J]=a$ $[aa,J]=2aa$ $[aaa,J]=3aaa$ $[aab,J]=4aab$ $[abb,J]=5abb$
$a^{\dagger},J]=-a^{\dagger} [$$b$,J]=4bb[$b$,J]=4bb $aa^{\dagger},J]=aaa^{\dagger} [aa$$ab^{\dagger},J]=0$ $abb^{\dagger},J]=abb^{\dagger}$
$[b,J]=2b$ $[ab,J]=3ab$ $a^{\dagger}a,J]=aa^{\dagger}a [aa$$^{\dagger}a^{\dagger}b,J]=0$ $ab^{\dagger}b,J]=ab^{\dagger}b$
$b^{\dagger},J]=-2b^{\dagger} [$$^{\dagger}b,J]=a^{\dagger}b [aa$$^{\dagger}aa,J]=a^{\dagger}aa$ $aa^{\dagger}b,J]=2aa^{\dagger}b=2a^{\dagger}ab+2b [$$a^{\dagger}bb,J]=3a^{\dagger}bb$
$b^{\dagger},J]=-ab^{\dagger} [aa$$^{\dagger}a^{\dagger},J]=-aa^{\dagger}a^{\dagger} [aa$$^{\dagger}ab,J]=2a^{\dagger}ab$ $ab^{\dagger}b^{\dagger},J]=-3ab^{\dagger}b^{\dagger}$
$^{\dagger}b^{\dagger},J]=-3a^{\dagger}b^{\dagger} [aa$$^{\dagger}aa^{\dagger},J]=-a^{\dagger}aa^{\dagger}$ $aa^{\dagger}b^{\dagger},J]=-2aa^{\dagger}b^{\dagger}=-2a^{\dagger}ab^{\dagger}-2b^{\dagger} [$$a^{\dagger}bb^{\dagger},J]=-a^{\dagger}bb^{\dagger}$
$^{\dagger}a^{\dagger},J]=-2a^{\dagger}a^{\dagger} [aa$$^{\dagger}a^{\dagger}a,J]=-a^{\dagger}a^{\dagger}a$ $a^{\dagger}ab^{\dagger},J]=-2a^{\dagger}ab^{\dagger} [$$a^{\dagger}b^{\dagger}b,J]=-a^{\dagger}b^{\dagger}b$
$b^{\dagger}b^{\dagger},J]=-4b^{\dagger}b^{\dagger} [$$^{\dagger}a^{\dagger}a^{\dagger},J]=-3a^{\dagger}a^{\dagger}a^{\dagger} [aa$$^{\dagger}a^{\dagger}b^{\dagger},J]=-4a^{\dagger}a^{\dagger}b^{\dagger}$ $a^{\dagger}b^{\dagger}b^{\dagger},J]=-5a^{\dagger}b^{\dagger}b^{\dagger}$
Note that commutators with triple $b$’s will have the same values as those with triple $a$’s with an extra factor of 2. The commutator defines a linear transformation T(_)=[_,J]T(\ \_\ )=[\ \_\ ,J] from Hermitian to anti-Hermitian operators. By inspection, we can see the kernel of $T$ is given by $H_{ab}$ as defined in the main text. Energy conservation is sufficient to single out the frequency doubling Hamiltonian from the structure of eq. (4).
Appendix B Frequency doubling eqs.
@@
$\dot{\mathcal{O}}=i[H,\mathcal{O}]+\sum_{\nu}\left(L_{\nu}^{\dagger}\mathcal{O}L_{\nu}-\frac{1}{2}L_{\nu}^{\dagger}L_{\nu}\mathcal{O}-\frac{1}{2}\mathcal{O}L_{\nu}^{\dagger}L_{\nu}\right)\ .$ (35)
@@
$H_{a}=\omega a^{\dagger}a\ ,H_{b}=2\omega b^{\dagger}b\ ,H_{I}=\kappa\left(b^{\dagger}a^{2}+ba^{\dagger 2}\right)$ (37)
@@
$L_{a}=\sqrt{2\gamma_{a}}a\ ,L_{b}=\sqrt{2\gamma_{b}}b\ .$ (38)
@@
$\alpha=ae^{i\omega t}\ ,\beta=be^{2i\omega t}$ (39)
Note that $\alpha$ and $\beta$ have the same commutation relations as $a,b$. Eq. (35) for modes $a$ and $b$ then reads,
$\dot{\alpha}$ $=$ $-2i\kappa\beta\alpha^{\dagger}-\gamma_{a}\alpha$ (40)
$\dot{\beta}$ $=$ $-i\kappa\alpha^{2}-\gamma_{b}\beta$ (41)
@@ C.1 Hamiltonian
$H_{O}=-\dfrac{\omega_{c}}{4}c^{\dagger}c\int\dfrac{d\bm{k}}{\sqrt{(2\pi)^{3}}}\left(\sqrt{\dfrac{8\pi G}{k}}\mathfrak{b}_{\bm{k}}+h.c.\right)\ ,$ (43)
where $k=|\bm{k}|=\omega_{k}$ is the GW frequency for the mode $\bm{k}$, the operator $\mathfrak{b}_{k}^{\lambda}$ ($\mathfrak{b}_{\bm{k}}^{\lambda\dagger}$) is the canonical graviton annihilation (creation) operator and $\omega_{c}$ is the cavity frequency with photon annihilation (creation) operator $c\ (c^{\dagger})$. To obtain a discrete version of (43) we introduce a quantization volume $V$ and note that $\left[\sqrt{8\pi G/k}\right]=L^{3/2}$, where $L$ denotes dimension of length. Moreover, $\left[d\bm{k}\right]=L^{-3}$. The graviton annihilation and creation operators then have dimension $\left[\mathfrak{b}_{\bm{k}}\right]=L^{3/2}$. Define $\mathfrak{b}_{\bm{k}}=\sqrt{V}\bm{b}_{\bm{k}}$, where $\bm{b}_{\bm{k}}$ is dimensionless. The discrete limit $d\bm{k}\rightarrow 1/V$, $(2\pi)^{-3/2}\int\rightarrow\sum$ leads to,
$H_{O}=-\dfrac{\omega_{c}}{4}c^{\dagger}c\sum_{\bm{k}}\left(\sqrt{\dfrac{8\pi G}{kV}}\bm{b}_{\bm{k}}+h.c.\right).$ (44)
@@
$H_{F}=\omega_{c}c^{\dagger}c+\sum_{\bm{k}}\omega_{k}\bm{b}_{\bm{k}}^{\dagger}\bm{b}_{\bm{k}}\,.$ (45)
@@
$U_{\bm{k}}(t)=e^{iB_{k}(t)(c^{\dagger}c)^{2}}e^{q_{\bm{k}}c^{\dagger}c(\gamma_{k}\bm{b}^{\dagger}_{k}-\gamma_{k}^{*}\bm{b}_{k})}\,$, (47)
@@
$\gamma_{k}$ $=$ $(1-e^{-i\omega_{k}t})$ (48)
$B_{k}(t)$ $=$ $q_{k}^{2}\left(\omega_{k}t-\sin\omega_{k}t\right)$ (49)
@@
$|\Psi(t)\rangle=\prod_{\bm{k}}e^{-i\bm{b}_{\bm{k}}^{\dagger}\bm{b}_{\bm{k}}\omega_{\bm{k}}t}|\Psi\rangle\,.$ (50)

### v2

@@ 6 Ringdown nonlinearities as detectors
In quantum optics, it has long been recognized that the output of a nonlinear process strongly depends on the particle number statistics of the input parent modes [38, 39, 40]. For the case of SHG, this effect can be understood through simple physical reasoning. Particle number statistics determines whether the parent fields are bunched or anti-bunched [25]. Energy conservation then requires the presence of two or more particles in mode $\omega$ for the creation of a particle in mode $2\omega$. Hence, bunched states are converted more efficiently than anti-bunched ones in SHG [42]. By analogy to quantum optics, we can therefore anticipate that the properties of nonlinearly generated GWs will depend on the graviton number statistics and other higher-order moments of the parent modes. Given the strongly nonlinear character of gravity, these effects can be significant in ringdowns.
@@
To quantify how strong nonlinear GW couplings in ringdowns can be, we may estimate the “graviton conversion efficiency” between modes, The mean number of gravitons in a GW of frequency $\omega$ and amplitude $A_{\omega}$ is [13],
$\langle N_{\omega}\rangle\approx\frac{V}{32\pi G}\omega A_{\omega}^{2}$ (6.24)
@@
$\eta_{\mathrm{ringdown}}=\frac{\langle N_{2\omega}\rangle}{\langle N_{\omega}\rangle}\approx 2\mathcal{R}A_{2\omega}$ (6.25)
where we recall the gravitational wave amplitude $A_{2\omega}\equiv A_{(4,4)}^{(2,2,0)\times(2,2,0)}$ discussed in section 2. At the source, GW amplitudes can be of order unity [81]. Taking $A^{(2,2,0)\times(2,2,0)}_{(4,4)}\sim 0.1$ and $\mathcal{R}\sim 0.15$ [29] we have the conversion efficiency,
$\eta_{\mathrm{ringdown}}\sim 10^{-2}\ .$ (6.26)
@@
$\eta_{\mathrm{gertsenshtein}}\sim 10^{-33}$ (6.27)
while for bar detectors [18],
$\eta_{\mathrm{bar}}\sim 10^{-36}$ (6.28)
This comparison shows that ringing BHs are essentially strongly coupled nonlinear detectors, with a high conversion efficiency between different modes. Nonlinear ringdowns may therefore be used as efficient detectors for quantum properties of GWs.
The graviton conversion efficiency (6.26) can also be used to estimate the departure from classicality during a nonlinear ringdown. To do that, let us consider Mandel’s $\mathcal{Q}$ parameter, defined as
$\mathcal{Q}\equiv\frac{\langle\Delta N^{2}\rangle-\langle N\rangle}{\langle N\rangle}\ .$ (6.29)
Classical states have $\mathcal{Q}\geq 0$, while $\mathcal{Q}<0$ implies sub-Poissonian statistics, which cannot be described classically [83]; $\mathcal{Q}$ can therefore be used to distinguish classical from quantum radiation. During SHG, an initial coherent state $|\alpha_{\omega}\rangle$ in the fundamental mode $\omega$ evolves over short times to a non-classical state with [69]
$\mathcal{Q}\sim-2\kappa^{2}t^{2}\ |\alpha_{\omega}|^{2}+\mathcal{O}(\kappa^{4}t^{4})$ (6.30)
where this expression is valid for $\kappa t|\alpha_{\omega}|\ll 1$. With the help of eq. (3.5), $\kappa t\sim|\alpha_{2\omega}|/|\alpha_{\omega}|^{2}$, hence
$\mathcal{Q}\sim-2|\alpha_{2\omega}|^{2}/|\alpha_{\omega}|^{2}\sim-\eta_{\rm ringdown}\ ,$ (6.31)
i.e. at short interaction times the fundamental mode will develop a small, but non-negligible negative Mandel parameter on the order of $\mathcal{Q}\sim-0.01$. The fundamental mode becomes non-classical. Note second harmonic also generates squeezing in the fundamental mode [69].
Let us now see an example illustrating how the statistical properties of parent states can become imprinted upon nonlinear modes in the case of SHG. Consider the dynamical eqs. (4.6) and (4.7). We take mode $a$ to be initially populated by some quantum state $\rho_{a}$, while $b$ is initially in the vacuum state. For short times, the mean number of particles in mode $b$ grows as [39],
$\langle N_{b}(t)\rangle=g^{(2)}(0)\ \langle N_{a}(0)\rangle^{2}\ \kappa^{2}t^{2},$ (6.32)
@@
$g^{(2)}(0)=\frac{\langle a^{\dagger}a^{\dagger}aa\rangle}{\langle a^{\dagger}a\rangle^{2}}$ (6.33)
is the second-order field auto-correlation function and expectation values are taken with respect to the initial state $\rho_{a}$. In deriving (6.32) we assumed $\sqrt{2g^{(2)}(0)\langle N_{a}(0)\rangle}\ \kappa t\ll 1$ and $t\ll\gamma^{-1}$, but note $\langle N_{b}(t)\rangle$ also exhibits significant dependence on the parent quantum state beyond the short time or small damping approximation [39].
The result (6.32) tells us that for short times, the efficiency with which particles are converted from modes $a$ to $b$ is directly proportional to the field auto-correlation function (6.33), which is related to the bunching or anti-bunching character of the wave [25]. For coherent states $g^{(2)}(0)=1$, for thermal states $g^{(2)}(0)=2$ and squeezed states have $g^{(2)}(0)=3+\frac{1}{\langle N_{a}(0)\rangle}$. Importantly, for large mean number of gravitons, squeezed states have $g^{(2)}(0)\approx 3$, independent of $\langle N_{a}(0)\rangle$. Therefore, we can expect second harmonic GW generation in BH ringdowns to be significantly enhanced for parent thermal and squeezed states in comparison to coherent states.
@@
$\langle Q_{b}(t)\rangle\approx\frac{\kappa t}{2}\Bigl<Q_{a}(0)P_{a}(0)+P_{a}(0)Q_{a}(0)\Bigr$> (6.34)
which to leading order remains valid in the presence of dissipation; see D for details.
@@
$\frac{1}{2}\Bigl<Q_{a}P_{a}+P_{a}Q_{a}\Bigr>=\langle Q_{a}\rangle\langle P_{a}\rangle+(2\bar{n}+1)\sinh r\sin\phi\cos\phi$ (6.35)
The first term on the r.h.s. is due to the coherent displacement of the GW and is the only term present for coherent states. In contrast, the second term depends on non-coherent properties of the state, the mean thermal occupation number $\bar{n}$ and squeezing parameter $\xi$. Note the non-coherent term is accompanied by an exponential enhancement characteristic of squeezing [13, 17].
As we have seen, interferometric detectors are sensitive to the mean GW quadratures [see (5.23)]. Eq. (6.34) is interesting, as it shows that SHG encodes information on the covariance of $a$ in the mean quadrature of $b$, which can be read-out by interferometric detectors. Detection of $\langle Q_{b}(t)\rangle$ by standard GW observations could therefore signal deviations of mode $a$ from a coherent state. For a discussion on the detectability of quadratic QNMs see [85].
7 Discussion
In conclusion, SHG has been observed in numerical simulations of nonlinear BH ringdowns [29, 30]. In quantum mechanics, second harmonics can produce non-classical states both in the parent and nonlinear modes, starting from initial coherent states [69]. This can be understood in terms of particle creation and annihilation: nonlinearities affect particle number distributions, leading to states which classical theory cannot account for [26].
@@ Appendix A Energy conservation and frequency doubling
$[a,a^{\dagger}]=[b^{\dagger},b]=\mathds{1}$ (A.36)
$[a,b]=[a^{\dagger},b]=[a,b^{\dagger}]=[a^{\dagger},b^{\dagger}]=0$ (A.37)
@@
$J=N_{a}+2N_{b}$ (A.38)
For convenience we list all the commutators with third-order operators in Table 1.
Table 1: Commutators with $J$.
$[a,J]=a$ $[aa,J]=2aa$ $[aaa,J]=3aaa$
$a^{\dagger},J]=-a^{\dagger} [$$b$,J]=4bb[$b$,J]=4bb $aaa^{\dagger},J]=aaa^{\dagger}$
$[b,J]=2b$ $[ab,J]=3ab$ $aa^{\dagger}a,J]=aa^{\dagger}a$
$b^{\dagger},J]=-2b^{\dagger} [$$^{\dagger}b,J]=a^{\dagger}b [aa$$^{\dagger}aa,J]=a^{\dagger}aa$
$b^{\dagger},J]=-ab^{\dagger} [aa$$a^{\dagger}a^{\dagger},J]=-aa^{\dagger}a^{\dagger}$
$^{\dagger}b^{\dagger},J]=-3a^{\dagger}b^{\dagger} [aa$$^{\dagger}aa^{\dagger},J]=-a^{\dagger}aa^{\dagger}$
$^{\dagger}a^{\dagger},J]=-2a^{\dagger}a^{\dagger} [aa$$^{\dagger}a^{\dagger}a,J]=-a^{\dagger}a^{\dagger}a$
$b^{\dagger}b^{\dagger},J]=-4b^{\dagger}b^{\dagger} [$$a^{\dagger}a^{\dagger}a^{\dagger},J]=-3a^{\dagger}a^{\dagger}a^{\dagger}$
$[aab,J]=4aab$ $[abb,J]=5abb$
$aab^{\dagger},J]=0 [$$abb^{\dagger},J]=abb^{\dagger}$
$a^{\dagger}a^{\dagger}b,J]=0 [$$ab^{\dagger}b,J]=ab^{\dagger}b$
$aa^{\dagger}b,J]=2aa^{\dagger}b=2a^{\dagger}ab+2b [$$a^{\dagger}bb,J]=3a^{\dagger}bb$
$a^{\dagger}ab,J]=2a^{\dagger}ab [$$ab^{\dagger}b^{\dagger},J]=-3ab^{\dagger}b^{\dagger}$
$aa^{\dagger}b^{\dagger},J]=-2aa^{\dagger}b^{\dagger}=-2a^{\dagger}ab^{\dagger}-2b^{\dagger} [$$a^{\dagger}bb^{\dagger},J]=-a^{\dagger}bb^{\dagger}$
$a^{\dagger}ab^{\dagger},J]=-2a^{\dagger}ab^{\dagger} [$$a^{\dagger}b^{\dagger}b,J]=-a^{\dagger}b^{\dagger}b$
$a^{\dagger}a^{\dagger}b^{\dagger},J]=-4a^{\dagger}a^{\dagger}b^{\dagger} [$$a^{\dagger}b^{\dagger}b^{\dagger},J]=-5a^{\dagger}b^{\dagger}b^{\dagger}$
Note that commutators with triple $b$’s will have the same values as those with triple $a$’s with an extra factor of 2. The commutator defines a linear transformation T(_)=[_,J]T(\ \_\ )=[\ \_\ ,J] from Hermitian to anti-Hermitian operators. By inspection, we can see the kernel of $T$ is given by $H_{ab}$ as defined in the main text. Energy conservation is sufficient to single out the frequency doubling Hamiltonian from the structure of eq. (3.4).
Appendix B Frequency doubling equations
@@
$\dot{\mathcal{O}}=i[H,\mathcal{O}]+\sum_{\nu}\left(L_{\nu}^{\dagger}\mathcal{O}L_{\nu}-\frac{1}{2}L_{\nu}^{\dagger}L_{\nu}\mathcal{O}-\frac{1}{2}\mathcal{O}L_{\nu}^{\dagger}L_{\nu}\right)\ .$ (B.39)
@@
$H_{a}=\omega a^{\dagger}a\ ,H_{b}=2\omega b^{\dagger}b\ ,H_{I}=\kappa\left(b^{\dagger}a^{2}+ba^{\dagger 2}\right)$ (B.41)
@@
$L_{a}=\sqrt{2\gamma_{a}}a\ ,L_{b}=\sqrt{2\gamma_{b}}b\ .$ (B.42)
@@
$\alpha=ae^{i\omega t}\ ,\beta=be^{2i\omega t}$ (B.43)
Note that $\alpha$ and $\beta$ have the same commutation relations as $a,b$. Eq. (B.39) for modes $a$ and $b$ then reads,
$\dot{\alpha}=-2i\kappa\beta\alpha^{\dagger}-\gamma_{a}\alpha$ (B.44)
$\dot{\beta}=-i\kappa\alpha^{2}-\gamma_{b}\beta$ (B.45)
@@ C.1 Hamiltonian
$H_{O}=-\dfrac{\omega_{c}}{4}c^{\dagger}c\int\dfrac{d\bm{k}}{\sqrt{(2\pi)^{3}}}\left(\sqrt{\dfrac{8\pi G}{k}}\mathfrak{b}_{\bm{k}}+h.c.\right)\ ,$ (C.47)
where $k=|\bm{k}|=\omega_{k}$ is the GW frequency for the mode $\bm{k}$, the operator $\mathfrak{b}_{k}^{\lambda}$ ($\mathfrak{b}_{\bm{k}}^{\lambda\dagger}$) is the canonical graviton annihilation (creation) operator and $\omega_{c}$ is the cavity frequency with photon annihilation (creation) operator $c\ (c^{\dagger})$. To obtain a discrete version of (C.47) we introduce a quantization volume $V$ and note that $\left[\sqrt{8\pi G/k}\right]=L^{3/2}$, where $L$ denotes dimension of length. Moreover, $\left[d\bm{k}\right]=L^{-3}$. The graviton annihilation and creation operators then have dimension $\left[\mathfrak{b}_{\bm{k}}\right]=L^{3/2}$. Define $\mathfrak{b}_{\bm{k}}=\sqrt{V}\bm{b}_{\bm{k}}$, where $\bm{b}_{\bm{k}}$ is dimensionless. The discrete limit $d\bm{k}\rightarrow 1/V$, $(2\pi)^{-3/2}\int\rightarrow\sum$ leads to,
$H_{O}=-\dfrac{\omega_{c}}{4}c^{\dagger}c\sum_{\bm{k}}\left(\sqrt{\dfrac{8\pi G}{kV}}\bm{b}_{\bm{k}}+h.c.\right).$ (C.48)
@@
$H_{F}=\omega_{c}c^{\dagger}c+\sum_{\bm{k}}\omega_{k}\bm{b}_{\bm{k}}^{\dagger}\bm{b}_{\bm{k}}\,.$ (C.49)
@@
$U_{\bm{k}}(t)=e^{iB_{k}(t)(c^{\dagger}c)^{2}}e^{q_{\bm{k}}c^{\dagger}c(\gamma_{k}\bm{b}^{\dagger}_{k}-\gamma_{k}^{*}\bm{b}_{k})}\,$, (C.51)
@@
$\gamma_{k}=(1-e^{-i\omega_{k}t})$ (C.52)
$B_{k}(t)=q_{k}^{2}\left(\omega_{k}t-\sin\omega_{k}t\right)$ (C.53)
@@
$|\Psi(t)\rangle=\prod_{\bm{k}}e^{-i\bm{b}_{\bm{k}}^{\dagger}\bm{b}_{\bm{k}}\omega_{\bm{k}}t}|\Psi\rangle\,.$ (C.54)